40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Entry  Tue Sep 14 20:14:37 2010, tara, DailyProgress, PMC, PMC open loop TF pmc_bode.png
    Reply  Mon Oct 4 21:05:21 2010, tara, DailyProgress, PMC, PMC open loop TF 
Message ID: 371     Entry time: Mon Oct 4 21:05:21 2010     In reply to: 360
Author: tara 
Type: DailyProgress 
Category: PMC 
Subject: PMC open loop TF 

 I checked the PSL setup today, and the PMC gain setup has to be changed from 30 dB (maximum on gain slider) to 22.5 for best transmission signal.


From the previous  setup, see quote below, the maximum gain of 30 dB on the PMC gain slider was not high enough,

This means that even though the gain is set to maximum, the signal from the transmitted light does not oscillate.

But today it did oscillate, and I had to reduce the gain to 22.5 dB. When I checked the PMC gain, I turned off the FSS loop

to make sure that the FSS loop won't actuate on the NPRO, and the signal are purely from PMC loop.

There are no other changes of parameters. The power input is 30.7 mW, RF power, phase shift are the same

as before.

After adjusting the PMC's gain, I also roughly adjust FSS's loop gain. I haven't optimized it yet, just determined it by looking

at C3:PSL-RCAV_RCTRANSPD signal on the oscilloscope.

common gain is changed from 5 to 8 dB

Fast gain is changed from 8.5 to 13 dB.

I saved change the values for PMC gain, common gain and fast gain in STARTUP.CMD file.

I don't know what causes the gain changes here, I will check the TF of PMC loop again. If it is real, it means

we might not have to modify the PDH servo card.

NOTE: the slowDC for locking both cavities is ~ -0.103 V



Today I measured open loop transfer function of PMC loop.

The measurement has two parts.

First the swept sine signal is sent to FP2 test point, TP4 is connected to A, TP3 is connected to B,

the magnitude, B/A, gives us [C][D][E] .

For the second part, the swept sine is sent to ext DC channel,

TP3 is connected to A, and TP4 is connected to B.

this is the TF of [F][A][B]


 [A]--<FP2>-----[B]-----{TP4} ------[C]-----[D]---<Ext DC> ----[E] ---- {TP3}----[F]-----> back to [A]





The magnitudes and phase from both measurements are added up

to get the whole open loop TF of PMC loop [A][B][C][D][E][F].

UGF is ~1k Hz.

PMC setup

Gain 30dB (mzx)

LO PWR 0.585

Power input 30.9 mW

PMC_PHCON 2.5 + 180 flip


I'll verify that the schematic matches up with the real circuit we are using.


ELOG V3.1.3-